OpenStudy (anonymous):

Express the given quantity as a single logarithm. 1/9 ln(x+2)^9 + 1/2[lnx-ln(x^2+3x+2)^2]

OpenStudy (anonymous):

1/9 ln(x+2)^9=ln(x+2) because a*lnx=ln(a^x) 1/2 lnx -1/2*ln(x^2+3x+2)^2=ln(sqrtx)-ln(x^2+3x+2)